Grade 12Algebra

Are the following statements true?
45 km : 60 km = 12 hours : 15 hours

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

45 km : 60 km = 12 hours : 15 hours
45 / 60 = 3 / 4
12 / 15 = 4 / 5
Hence, False

Thanks